Verify
(1+tanx/sinx)-secx=cscx

Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

Verify
(1+tanx/sinx)-secx=cscx
1 + 1/cos - 1/cos = 1/sin
1 = 1/sin It's not an identity, not equal

Answer by lwsshak3(11628)

Verify
(1+tanx/sinx)-secx=cscx
..
Start from left-hand side
(1+sin/cos)/sin)-1/cos
(cos+sin/sincos)-1/cos
LCD=sincos
(cos+sin-sin)/sincos
cos/sincos=1/sin=csc
verified: left-hand side=right-hand side